1.    DBS Farms is a producer and retailer of farm products. DBS main products are Mangoes, Pawpaw and Pineapples. The current price of the Mangoes per Kilogram is GHS 50, the Pawpaw/Kg is GHS 80 and the Pineapple is GHS 40. This year the DBS Farms sold 10,000 kgs of Mangoes, 20,000 kgs of Pawpaw and 1 million kgs of Pineapples. In an attempt to improve revenue, the managers of the firm have decided to increase all prices by 10%. Market research has suggested that the price elasticity of demand for each product is: Mangoes: - 1.5; Pawpaw: -2.5; Pineapples: - 0.6. You have been asked to evaluate the planned price increases.

 

a.      Comment on the planned price changes.

  1. Would a 10% price reduction have been better for some or all of the products?
1
Expert's answer
2021-05-04T07:20:32-0400

Given 

The current price of the Mangoes per Kilogram is GHS 50 and the quantity is 10,000 kgs of mangoes. 

The current price of the Pawpaw/Kg is GHS 80 and the quantity is 20,000 kgs of Pawpaw. 

The current price of the Pineapple is GHS 40 and the quantity is 1 million kgs of pineapples. 

The price elasticity of demand for each product is:

Mangoes: - 1.5

Pawpaw: -2.5

Pineapples: - 0.6.


Total revenue in the current prices are:

Total revenue from Mangoes "= 50 \u00d7 10,000 = 500,000."

Total revenue from Pawpaw "= 80 \u00d7 20,000 = 1,600,000."

Total revenue from pineapples "= 40 \u00d7 1 million = 40 million."

Total revenue from all fruits "= 500,000 + 1,600,000 + 40,000,000 = 42,100,000."


Now the new prices and the quantity. 

Price elasticity of demand "=\frac{\%\Delta Q}{\%\Delta P}"


For Mangoes. 

New price = 55. New quantity = 85,000 

New total revenue = 55 × 85000 = 467,500. 


For Pawpaw. 

New price = 88, New quantity = 15,000. 

New total revenue = 88 × 15,000 = 1,320,000 


For pineapples. 

New price = 44, New quantity = 9,400,000. 

Total revenue = 41,360,000.

 

New total revenue from all the fruits "= 467,500 + 1,320,000 + 41,360,000 = 43,147,500."


price increse by 10% has given increase in revenue but it is advisable to increse price only in pineapple.

price reduction 10% will be advisable in mango and pawpaw based on price elasticty of demand.